# Data Selection via Optimal Control for Language Models

### v1

@@ Appendix A Connection Between AUC and Scaling Law Constants
We show that the area under the loss curve (AUC) is directly connected to the scaling law [44] constants, and minimizing AUC essentially improves the scaling properties of LMs. As suggested by existing literature [44, 40], the test losses of LMs follow a power-law with respect to the training steps $t$ after a warmup stage:
@@
To prove Theorem 2.1, we first describe the standard discrete-time Pontryagin’s Maximum Principle (PMP; 63) in Optimal Control [47] for time-variant control variables:
@@
Proof of Theorem B.1 can be found in various textbooks in Optimal Control [47, 23]. In our context, we interpret ${\bm{\theta}}_{t}$ as the model parameters, $J(\cdot)$ as the downstream loss function, $U$ as the $|\mathcal{D}|$-dimensional simplex as defined in Section 2.1 and $f$ as the GD operation where the data weights ${\bm{\gamma}}_{t}$ changes with respect to the training steps $t$:
@@ Appendix C Derivation for Adam
We provide our formulation and derivation for Adam [45] in this section. For $10\leq t\leq T-$ 1, the parameter update rules of Adam is given by
@@
where $\beta_{1}$, $\beta_{2}$, $\epsilon$, $\eta$ are hyper-parameters and ${\bm{g}}_{t}=\nabla L({\bm{\theta}}_{t},{\bm{\gamma}}_{t},\mathcal{D})$. We set ${\bm{m}}_{0}=\bm{0}$ and ${\bm{v}}_{0}=\bm{0}$. To formulate the LM training with Adam as an Optimal Control [47] problem, we treat the vector $\bm{\Theta}_{t}=\left[{\bm{\theta}}_{t},{\bm{m}}_{t},{\bm{v}}_{t}\right]^{\top}\in\mathbb{R}^{3N}$ as the state variable. Let $F$ denote the state-transition function from $\bm{\Theta}_{t}$ to $\bm{\Theta}_{t+1}$:
@@
Similar to GD, we can still define the Hamiltonian function of the problem and obtain the following theorem with Pontryagin’s Maximum Principle (PMP; 63):
@@ Appendix D Algorithm 1 as Proximal Gradient Decent
We provide another view of Algorithm 1 as using Proximal Gradient Decent method [6]. Specifically, we can optimize Eq. (3) with the following rules:
@@
Appendix E Implementation Details
E.1 Solving Data Quality Scores
@@
Algorithm 1 needs to compute the Hessian matrix, as in Eq. (5), and per-instance vector-gradient product, as in Eq. (6). We adopt the Jacobian-Vector-Product [^4] https://pytorch.org/docs/stable/func.api.html in PyTorch [62] to efficiently implement these operations. There is still a large room for this implementation, such as adopting other deep learning frameworks [19, 11] or using lower-complexity algorithms [75]. Algorithm 1 requires storing all the LM checkpoints ${\bm{\theta}}_{t}$ from $t=0$ to $t=T-1$ in the forward inner loop for computing ${\bm{\lambda}}_{t}$ in the reverse inner loop. To reduce the single-device GPU memory use, we implement an activation partition algorithm inspired by ZeRO-2 [64], where ${\bm{\theta}}_{t}$ in one inner-loop pass are stored in different GPU devices. We also adopt a strategy inspired by activation checkpointing [14].
E.2 Training Data Scorer
@@
We fine-tune the Fairseq-Dense-125M model [5] on the solved data weights ${\bm{\gamma}}^{*}$ to obtain the data scorer. as in Eq. (8), we adopt a linear transformation on the mean pooling of the instance’s representations along the sequence length. The size of the hidden state is 768. We optimize Eq. (8) with the AdamW [54] optimizer for 5 epochs, using a learning rate $1\times 10^{-4}$ and a batch size 512. We split 10% samples from $\mathcal{D}^{\mathrm{prx}}$ for validation and select the checkpoint achieving the highest Spearman correlation score on the validation set to infer data quality scores in $\mathcal{D}$.
E.3 Pre-Training
@@
We pre-train all our models for about 50B tokens with a batch size of 512 and a max input length of 1,024. We use the AdamW [54] optimizer and cosine learning rate scheduler, which warmups the learning rates for 2K steps and then decays it to 10% of the maximal value. We follow [12] to set the maximal learning rates as listed in Table 6, together with the model configurations.
@@
E.4 Evaluation
We evaluate our trained models on MMLU [39] and the evaluation sets used in OLMo [31], including Hellaswag (HS; 82), Winograde (Wino.; 46), LAMBADA (LAMB; 61), OpenbookQA (OBQA; 57), ARC-easy/challenge (ARC-e/c; 17), PIQA [9], SciQ [76], and BoolQ [16]. We adopt the LM-evaluation-harness library [27] [^5] https://github.com/t1101675/lm-harness/ to conduct the zero-shot evaluation, which formulates the tasks as the multiple-choice problems and chooses the answer by comparing the answer-length-normed loss across candidates (acc_norm). We also compute the test loss of our trained models on the DCLM corpus [48], a 10K subset uniformly sampled from the high-quality pre-training corpus in Li et al. [48].
E.5 Baselines
Reducible Holdout Loss Selection (RHO-Loss; 58, 51).
@@ E.5 Baselines
Data Selection via Importance Resampling (DSIR; 79).
DSIR selects pre-training data based on the n-gram feature overlap between the instances in the downstream dataset (LIMA [84] in our experiments) and the large-scale corpus. Pre-training instances whose features have high probabilities under the feature distribution of the downstream dataset will obtain higher sampling weights. We use the official implementation of DSIR [^6] https://github.com/p-lambda/dsir.
Influence Score (IF-Score 43, 32).
We adopt the influence function [43, 32] to measure the quality of $x$ using the downstream loss $J({\bm{\theta}})$ as the target:
@@
where ${\bm{\theta}}_{T}$ is the LM parameters trained for $T$ steps and $L({\bm{\theta}}_{T})=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}l(x,{\bm{\theta}}_{T})$. We adopt a linear-time iterative algorithm to compute the inverse-Hessian-vector-product [2]. To reduce the computational cost, we adopt a similar efficient implementation as PDS by computing the IF-Scores on a small proxy data based on a small proxy LM and then transferring these scores to a large pre-training corpus with a fine-tuned data scorer. We select examples with the top 40% scores inferred by the data scorer on the large pre-training corpus.
E.6 Simulated Setting for Experiments in Table 6
To exactly run Algorithm 1 with a feasible computational overhead, we adopt a 12M LM from the Mistral [42] family, with a small vocabulary. We uniformly sample 4,096 instances as $\mathcal{D}$, with a max sequence length of 256, and construct a 16K vocabulary from $\mathcal{D}$ and LIMA. We run the inner loops for 5K steps and the outer loop for 5 epochs to get the PDS (exact) line in Figure 6. For PDS (Efficient), we adopt an 8.7M model as the proxy LM and set $M=5$, $T^{\mathrm{prx}}=100$. We run the inner loop using SGD with a batch size of 128. The outer loop epoch number is set to 1.
Appendix F Complexity Analysis
Following Hoffmann et al. [40], for an LM with $N$ parameters to be trained on $D$ tokens, we assume the computational FLOPs of a forward and a backward pass are $2ND$ and $4ND$, respectively. We compute the FLOPs and asymptotic complexities of different stages in PDS as follows:
@@ Appendix F Complexity Analysis
Appendix G More Results
G.1 Scaling Curves of Computation for Other Model Sizes
@@
Figure 8: Scaling curves of average accuracy on the OLMo [31] evaluation datasets with respect to computation for 160M, 470M, and 1B sizes.
@@
G.2 Test Loss Extrapolation with the Scaling Law

### v2

@@ 5 Conclusion
Acknowledgements
This work is supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604) and Tsinghua University Initiative Scientific Research Program.
@@ Appendix A Connection Between AUC and Scaling Law Constants
We show that the area under the loss curve (AUC) is directly connected to the scaling law (Kaplan et al., 2020) constants, and minimizing AUC essentially improves the scaling properties of LMs. As suggested by existing literature (Kaplan et al., 2020; Hoffmann et al., 2022), the test losses of LMs follow a power-law with respect to the training steps $t$ after a warmup stage:
@@
To prove Theorem 2.1, we first describe the standard discrete-time Pontryagin’s Maximum Principle (PMP; Pontryagin, 2018) in Optimal Control (Lewis et al., 2012) for time-variant control variables:
@@
Proof of Theorem B.1 can be found in various textbooks in Optimal Control (Lewis et al., 2012; Evans, 2005). In our context, we interpret ${\bm{\theta}}_{t}$ as the model parameters, $J(\cdot)$ as the downstream loss function, $U$ as the $|\mathcal{D}|$-dimensional simplex as defined in Section 2.1 and $f$ as the GD operation where the data weights ${\bm{\gamma}}_{t}$ changes with respect to the training steps $t$:
@@ Appendix C Derivation for Adam
We provide our formulation and derivation for Adam (Kingma & Ba, 2015) in this section. For $10\leq t\leq T-$ 1, the parameter update rules of Adam is given by
@@
where $\beta_{1}$, $\beta_{2}$, $\epsilon$, $\eta$ are hyper-parameters and ${\bm{g}}_{t}=\nabla L({\bm{\theta}}_{t},{\bm{\gamma}}_{t},\mathcal{D})$. We set ${\bm{m}}_{0}=\bm{0}$ and ${\bm{v}}_{0}=\bm{0}$. To formulate the LM training with Adam as an Optimal Control (Lewis et al., 2012) problem, we treat the vector $\bm{\Theta}_{t}=\left[{\bm{\theta}}_{t},{\bm{m}}_{t},{\bm{v}}_{t}\right]^{\top}\in\mathbb{R}^{3N}$ as the state variable. Let $F$ denote the state-transition function from $\bm{\Theta}_{t}$ to $\bm{\Theta}_{t+1}$:
@@
Similar to GD, we can still define the Hamiltonian function of the problem and obtain the following theorem with Pontryagin’s Maximum Principle (PMP; Pontryagin, 2018):
@@ Appendix D Algorithm 1 as Proximal Gradient Decent
We provide another view of Algorithm 1 as using Proximal Gradient Decent method (Bauschke & Combettes, 2011). Specifically, we can optimize Eq. (3) with the following rules:
@@
Appendix E Extension to Considering Data Dependence
In Section 2.1, we formulate our problem by focusing on the individual importance of each data point to the downstream loss. This potentially affects the performance of PDS when the original corpus to select from contain too much duplication. The reason is that, similar examples tend to have similar distances to ${\bm{\lambda}}$, which means that if one example individually get high $\gamma$, the others will also get high $\gamma$ and be selected, affecting the diversity of the selected corpus. In this section, we provide a discussion on how the dependence between the selection data points can be considered to improve the diversity of the selected data, as an extension to our theoretical framework. Specifically, we can improve the pair-wise difference between the examples to promote the diversity of the selected data by considering the following pair-wise similarity:
$\operatorname{sim}(x_{n},x_{m})=\frac{\bm{h}_{n}^{\top}\bm{h}_{m}}{||\bm{h}_{n}||||\bm{h}_{m}||},$ (51)
where $\bm{h}_{n}$ and $\bm{h}_{m}$ are the representations of each example generated by a model like RoBERTa (Liu et al., 2019). Then, we have the following loss to minimize, which encourages the difference between examples:
$\mathcal{L}^{\text{diversity}}=\sum_{n,m}\text{sim}(x_{n},x_{m})\gamma_{n}\gamma_{m}=\bm{\gamma}^{\top}\bm{S}\bm{\gamma},$ (52)
where $\bm{S}=\left[\operatorname{sim}(x_{n},x_{m})\right]_{1\leq n\leq|\mathcal{D}|,1\leq m\leq|\mathcal{D}|}$. We can add this loss to the optimization problem in Eq. (3), using a hyper-parameter $w$ to control its weight. Additionally, is it also useful to add an $l_{0}$-norm constraint on $\gamma$ to constrain the number of non-zero elements to the selected data number $K$. This is because when pairwise interactions between examples are considered, adding or removing one data point would affect the contribution of the others. Therefore, the optimization problem becomes:
$\min_{\bm{\gamma}}$ $\sum_{t=1}^{T}J({\bm{\theta}}_{t})+w\cdot\bm{\gamma}\bm{S}\bm{\gamma}$ (53)
s.t. ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta\nabla L({\bm{\theta}}_{t},{\bm{\gamma}}),\ \ {\bm{\gamma}}\in U,\ \ \sum_{n=1}^{N}\mathbbm{1}[\gamma_{n}>0]=K.$
However, the problem with the $l_{0}$-norm constraint is computationally intractable to solve because computing the $l_{0}$-norm involves determining the exact sparsity pattern, which is equivalent to solving a combinatorial optimization problem. For high-dimensional ${\bm{\gamma}}$, this becomes computationally prohibitive as the complexity grows exponentially with the size of the problem. Furthermore, it is unclear how the constraint would affect the physical property of the dynamics and resulting in the optimization problem. For example, whether it will introduce any singularity during the dynamic process or the inaccuracy from the approximated optimization procedure (given $l_{0}$-norm is not differentiable) would lead to any unexpected outcome. We leave the design of effective and efficient methods to solve the problem in Eq. (53) to future work.
Appendix F Discussion on Eq. (6) of Theorem 2.1
The Optimality.
In Theorem 2.1, Eq. (6) takes a maximum operation to get ${\bm{\gamma}}^{*}$, which increases the $\gamma$ scores with the highest $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta^{*})$ values and reduce those $\gamma$ with lower $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta^{*})$ values to 0. However, this does not imply that a single training example is the optimal for data selection for the following reasons:
  • First, there could exist several samples having the same largest $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta^{*})$ values when the model’s parameter $\theta^{*}$ is trained under the optimal ${\bm{\gamma}}$ scores, i.e. $\exists S=\{n_{1},n_{2},\cdots,n_{|S|}\}$ such that for any $m\notin S$.
    $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n_{1}},\theta^{*})=\cdots=\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n_{|S|}},\theta^{*})>\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{m},\theta^{*})$ (54)
    Therefore, multiple samples in $S$ can have non-zero $\gamma^{*}$ scores to satisfy Eq. (6), as long as those samples not in $S$ get zero $\gamma^{*}$ scores. Accordingly, although “a single training sample” is a solution to Eq. (6), there also exist many more “equally qualified” solutions (such as assigning uniform $\gamma$ values to the data points in $S$) that satisfy Eq. (6).
  • Second, Theorem 2.1 is a necessary condition of the $\gamma$‘s optimality. This is because the two techniques we used to prove Theorem 2.1 in Appendix B, i.e., (1) time-variant PMP for discrete dynamical system (Theorem B.1) and (2) The Lagrange Multiplier method, are all necessary conditions. Therefore, the unreasonable “a single training sample” is not necessarily the optimal solution to the problem.
Many empirical results (Geering, 2007; Seierstad & Sydsaeter, 1986) have shown that solving this necessary condition can successfully lead to fairly good solutions, which is much like using gradient descent in deep learning even if it does not guarantee the optimality of the solution.
Intuition Behind ${\bm{\gamma}}$ Updates.
During the update of ${\bm{\gamma}}$ in Algorithm 1, since $\theta$ has not reached its optimum yet, the sample set $S$ having the same largest $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta)$ values has not emerged, making the max operation unstable in practice. Therefore, we increase $\gamma$ by a value proportional to $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta)$ in each iteration of the outer loop as in Algorithm 1. This updating strategy still satisfies our theory because it guarantees that the sample set having the largest $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta)$ values gets the highest $\gamma$ values. It also ensures that the samples with the same $\sum_{t=0}^{T}\lambda_{t+1}^{\top}\nabla l(x_{n},\theta)$ values receive the same $\gamma$ scores, which avoids the “single training sample” solution.
Appendix G Implementation Details
G.1 Solving Data Quality Scores
@@
Algorithm 1 needs to compute the Hessian matrix, as in Eq. (5), and per-instance vector-gradient product, as in Eq. (6). We adopt the Jacobian-Vector-Product [^4] https://pytorch.org/docs/stable/func.api.html in PyTorch (Paszke et al., 2019) to efficiently implement these operations. There is still a large room for this implementation, such as adopting other deep learning frameworks (Dagréou et al., 2024; Bradbury et al., 2018) or using lower-complexity algorithms (Wang et al., 2024). Algorithm 1 requires storing all the LM checkpoints ${\bm{\theta}}_{t}$ from $t=0$ to $t=T-1$ in the forward inner loop for computing ${\bm{\lambda}}_{t}$ in the reverse inner loop. To reduce the single-device GPU memory use, we implement an activation partition algorithm inspired by ZeRO-2 (Rajbhandari et al., 2020), where ${\bm{\theta}}_{t}$ in one inner-loop pass are stored in different GPU devices. We also adopt a strategy inspired by activation checkpointing (Chen et al., 2016).
G.2 Training Data Scorer
@@
We fine-tune the Fairseq-Dense-125M model (Artetxe et al., 2022) on the solved data weights ${\bm{\gamma}}^{*}$ to obtain the data scorer. as in Eq. (8), we adopt a linear transformation on the mean pooling of the instance’s representations along the sequence length. The size of the hidden state is 768. We optimize Eq. (8) with the AdamW (Loshchilov & Hutter, 2019) optimizer for 5 epochs, using a learning rate $1\times 10^{-4}$ and a batch size 512. We split 10% samples from $\mathcal{D}^{\mathrm{prx}}$ for validation and select the checkpoint achieving the highest Spearman correlation score on the validation set to infer data quality scores in $\mathcal{D}$.
G.3 Pre-Training
@@
We pre-train all our models for about 50B tokens with a batch size of 512 and a max input length of 1,024. We use the AdamW (Loshchilov & Hutter, 2019) optimizer and cosine learning rate scheduler, which warmups the learning rates for 2K steps and then decays it to 10% of the maximal value. We follow (Brown et al., 2020) to set the maximal learning rates as listed in Table 6, together with the model configurations.
@@
G.4 Evaluation
We evaluate our trained models on MMLU (Hendrycks et al., 2021) and the evaluation sets used in OLMo (Groeneveld et al., 2024), including Hellaswag (HS; Zellers et al., 2019), Winograde (Wino.; Levesque et al., 2012), LAMBADA (LAMB; Paperno et al., 2016), OpenbookQA (OBQA; Mihaylov et al., 2018), ARC-easy/challenge (ARC-e/c; Clark et al., 2018), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), and BoolQ (Clark et al., 2019). We adopt the LM-evaluation-harness library (Gao et al., 2024) [^5] https://github.com/t1101675/lm-harness/ to conduct the zero-shot evaluation, which formulates the tasks as the multiple-choice problems and chooses the answer by comparing the answer-length-normed loss across candidates (acc_norm). We also compute the test loss of our trained models on the DCLM corpus (Li et al., 2024), a 10K subset uniformly sampled from the high-quality pre-training corpus in Li et al. (2024).
G.5 Baselines
Reducible Holdout Loss Selection (RHO-Loss; Mindermann et al., 2022; Lin et al., 2024).
@@ G.5 Baselines
Data Selection via Importance Resampling (DSIR; Xie et al., 2023).
DSIR selects pre-training data based on the n-gram feature overlap between the instances in the downstream dataset (LIMA (Zhou et al., 2024) in our experiments) and the large-scale corpus. Pre-training instances whose features have high probabilities under the feature distribution of the downstream dataset will obtain higher sampling weights. We use the official implementation of DSIR [^6] https://github.com/p-lambda/dsir.
Influence Score (IF-Score Koh & Liang, 2017; Grosse et al., 2023).
We adopt the influence function (Koh & Liang, 2017; Grosse et al., 2023) to measure the quality of $x$ using the downstream loss $J({\bm{\theta}})$ as the target:
@@
where ${\bm{\theta}}_{T}$ is the LM parameters trained for $T$ steps and $L({\bm{\theta}}_{T})=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}l(x,{\bm{\theta}}_{T})$. We adopt a linear-time iterative algorithm to compute the inverse-Hessian-vector-product (Agarwal et al., 2017). To reduce the computational cost, we adopt a similar efficient implementation as PDS by computing the IF-Scores on a small proxy data based on a small proxy LM and then transferring these scores to a large pre-training corpus with a fine-tuned data scorer. We select examples with the top 40% scores inferred by the data scorer on the large pre-training corpus.
G.6 Simulated Setting for Experiments in Table 6
To exactly run Algorithm 1 with a feasible computational overhead, we adopt a 12M LM from the Mistral (Jiang et al., 2023) family, with a small vocabulary. We uniformly sample 4,096 instances as $\mathcal{D}$, with a max sequence length of 256, and construct a 16K vocabulary from $\mathcal{D}$ and LIMA. We run the inner loops for 5K steps and the outer loop for 5 epochs to get the PDS (exact) line in Figure 6. For PDS (Efficient), we adopt an 8.7M model as the proxy LM and set $M=5$, $T^{\mathrm{prx}}=100$. We run the inner loop using SGD with a batch size of 128. The outer loop epoch number is set to 1.
Appendix H Complexity Analysis
Following Hoffmann et al. (2022), for an LM with $N$ parameters to be trained on $D$ tokens, we assume the computational FLOPs of a forward and a backward pass are $2ND$ and $4ND$, respectively. We compute the FLOPs and asymptotic complexities of different stages in PDS as follows:
@@ Appendix H Complexity Analysis
Appendix I More Results
I.1 Results on 160M Models
We present the results on the OLMo (Groeneveld et al., 2024) evaluation sets based on the 160M models in Table 7. PDS achieves the best performance in most cases compared to the baselines.
HS LAMB Wino. OBQA ARC-e ARC-c PIQA SciQ BoolQ Avg.
Conventional 32.2 34.9 51.4 25.6 40.9 22.5 58.3 65.5 57.6 43.2
RHO-Loss 32.2 35.3 53.2 28.1 40.5 24.1 58.5 63.1 53.0 43.1
DSIR 32.8 35.7 52.5 26.6 41.2 23.8 57.8 68.7 54.7 43.8
IF-Score 32.2 35.7 51.1 27.4 40.8 22.6 57.6 64.1 55.8 43.0
PDS 33.5 38.2 51.4 28.4 42.3 24.1 59.2 68.8 58.7 45.0
Table 7: Downstream evaluation results on 160M models. We report the accuracy scores and the average scores across the datasets. The best scores of each model size are boldfaced. PDS achieves the best performance in most cases compared to the baselines.
I.2 Scaling Curves of Computation for Other Model Sizes
@@
Figure 8: Scaling curves of average accuracy on the OLMo (Groeneveld et al., 2024) evaluation datasets with respect to computation for 160M, 470M, and 1B sizes.
@@
I.3 Results on MMLU Measured by Perplexity on Ground Truth
As suggest by Schaeffer et al. (2023), using more smooth and continuous metrics like the LM’s perplexity on the ground truth better reveals the gap between different base models. Therefore, we include this results in Table 8 as a supplement to Table 2, which shows substantial improvement of PDS over the baselines.
Method PPL (160M) PPL (470M)
Conventional 44.7 34.8
RHO-Loss 42.1 33.0
DSIR 40.5 34.0
IF-Score 41.5 31.1
PDS 36.6 27.1
Table 8: The LMs’ perplexity (PPL) on the ground truth for zero-shot evaluation on MMLU (Hendrycks et al., 2021). The best scores of each model size are boldfaced.
I.4 Test Loss Extrapolation with the Scaling Law